OpenStudy (anonymous):

if a tugboat goes upstream 120 miles in 15 hours the return trip takes 10 hours find the speed of the boat and the speed of the current of the return trip? HELP PLEASE!

OpenStudy (anonymous):

upstream means opposite to flow of river so less total speed and more time downstream means with the flow of river and so more total speed and less time eq 120/(x-y)=15 and 120/(x+y)=10 solve

OpenStudy (anonymous):

what number gets plugged in for x-y i dont understand

OpenStudy (anonymous):

no number is plugged 2 different eq will be formed..solve the 2 simultaneous equations

OpenStudy (anonymous):

so you divide 120by x-y? How

OpenStudy (anonymous):

u can move 15 to denominator and x-y to right..similarly to other eq

hero (hero):

Use D = rt and r = (b ± c) to solve Upstream: 120 = (b - c)15 Downstream: 120 = (b + c)10

hero (hero):

Divide both sides of the first equation by 15; Divide both sides of the second equation by 10 to get: 8 = b - c 12 = b + c

hero (hero):

Now solve the system

OpenStudy (anonymous):

my problem is I don't understand how to get the b and c

hero (hero):

Isolate b in both equations, then set b = b

hero (hero):

If I isolate b in both equations I get b = 8 + c b = 12 - c Now setting b = b, I get 8 + c = 12 - c Can you finish solving for c?

OpenStudy (anonymous):

c=4

hero (hero):

8 + c = 12 - c c + c = 12 - 8 2c = 4 c = 4/2

OpenStudy (anonymous):

so the speed of the boat was 4 and the current is 2

hero (hero):

The current is 2, however, I don't agree with your speed of the boat

OpenStudy (anonymous):

omg I never going to understand this \

hero (hero):

Go to the equation where b was isolated: b = 8 + c Substitute c = 2 b = 8 + 2 b = ?

OpenStudy (anonymous):

10 how do you understand this stuff I just don't get it

hero (hero):

It's algebra. You just have to be familiar with the basic properties and formulas

OpenStudy (anonymous):

I need to you come sit with me at my house and be my teacher. LOL. thanks for your help

hero (hero):

For this problem, you need to understand the following 1. The relationship between distance, rate, and time (D = rt) 2. Upstream and Downstream 3. The relationship between current speed, boat speed and rate 4. Systems of Equations (Solving multiple equations, variable isolation and substitution)

hero (hero):

Would you say that you are familiar with all of the above?